40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 8 19:13:39 2019, Jon, Update, Upgrade, Ready for pumpdown tomorrow 
    Reply  Wed Jan 9 11:07:09 2019, gautam, Update, VAC, New Vac channel logging Screenshot_from_2019-01-09_11-08-28.png
Message ID: 14391     Entry time: Wed Jan 9 11:07:09 2019     In reply to: 14390
Author: gautam 
Type: Update 
Category: VAC 
Subject: New Vac channel logging 

Looks like I didn't restart all the daqd processes last night, so the data was not in fact being recorded to frames. I just restarted everything, and looks like the data for the last 3 minutes are being recorded yes. Is it reasonable that the TP1 current channel is reporting 0.75A of current draw now, when the pump is off? Also the temperature readback of TP3 seems a lot jumpier than that of TP2, probably has to do with the old controller having fewer ADC bits or something, but perhaps the SMOO needs to be adjusted.


Gautam and I updated the framebuilder config file, adding the newly-added channels to the list of those to be logged.

Attachment 1: Screenshot_from_2019-01-09_11-08-28.png  70 kB  | Hide | Hide all
ELOG V3.1.3-